OpenStudy (anonymous):

1+2+4+8+16+32.............. = ? Find the Sum

OpenStudy (anonymous):

64, 128,256,512

OpenStudy (anonymous):

1 (1+2+4+8+16+32.............. ) (2-1)(1+2+4+8+16+32.............. ) = 2 + 4 +8+ 16 +32 ....... -1-2-4-8-16-32.......... = -1

OpenStudy (anonymous):

its infinity

OpenStudy (anonymous):

Lol...i knw....Nw tell me wats the wrong in my sloution

OpenStudy (anonymous):

huh they haven't given the last term and btw its basically a geometric progression

OpenStudy (anonymous):

Do u Find any mistake in the way i did....

OpenStudy (anonymous):

yeah !!!!! there is a mistake in ur where did that -1 come from????

OpenStudy (anonymous):

2-1 = 1...................lol

OpenStudy (mayankdevnani):

limit help....

OpenStudy (mayankdevnani):

... 4, 8, 16, 32 ... 512 or 4(1), 4(2), 4(4), 4(8), ... , 4(128) or 4(2^0), 4(2^1), 4(2^2), 4(2^3), ... , 4(2^7) S(n) = 4 ∑ 2^k [ k = 0 to n ] ....... = 4 [ 2^(n+1) - 1 ] S(7) = 4 [ 2^(7+1) - 1 ] ....... = 1020

OpenStudy (anonymous):

Lol....i dont need the answer..) i want to knw is there any mistake in my solution

OpenStudy (mayankdevnani):

don't angry @Yahoo! it's my limit

OpenStudy (anonymous):

lol....i am not angry ...this is a Funny....question....

OpenStudy (mayankdevnani):

just apply my method and get the answer

OpenStudy (anonymous):

@Yahoo! : its not that let the series be A by ur terms => (2-1)A =>2A -A =A

OpenStudy (anonymous):

oho...Again..... Tell me is there any wrong...lol

OpenStudy (anonymous):

got it!!!!

OpenStudy (mayankdevnani):

\[\Huge{\bf{\color{blue}{i know you can }\color{red}{doit}}}\]

OpenStudy (anonymous):

I hate these! I can't see see what is wrong but obviously there is something as it should total infinity... I would just do the 2-1 in the bracket and get the same answer... (is it not that there will always be one last term in the positive set of numbers which is much bigger than its negative counterpart in that set, and eventually that is infinity..?)

OpenStudy (anonymous):

1 (1+2+4+8+16+32.............. ) (2-1)(1+2+4+8+16+32.............. ) = 2 + 4 +8+ 16 +32 ....... -1-2-4-8-16-32.......... = -1

OpenStudy (anonymous):

Also... first terms, 2-1 = 1, second terms: 4-2 = 2, so you DO get the original series back, and it does NOT equal -1.

OpenStudy (anonymous):

2+4+8+16+32..................+2n 1+2+4+8+16+32............. + n {subtracting} ----------------------------- 1+2+4+8+16 +32+...........n

OpenStudy (anonymous):

take make sense @DHASHNI

OpenStudy (anonymous):

WAT!!!

OpenStudy (anonymous):

*that

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

exactly @DHASHNI ! have a medal too :)

OpenStudy (anonymous):

You are subtracting sets of different cardinality. That's like adding unlike terms, you can't do it.

OpenStudy (unklerhaukus):

1 (1+2+4+8+16+32.............. ) (2-1)(1+2+4+8+16+32.............. ) = (2 + 4 +8+ 16 +32 +64+......)+( -1-2-4-8-16 -32 .........) = 63+.....